OpenStudy (anonymous):

A steel paper clip is bent by a force of 10N exerted on it through a distance of 5 cm. By how much does the temperature of the midsection of the paper clip increase if all the energy goes into the internal energy of this section, which has a mass of .1 g? Use 2100 J/kg°C for the specific heat capacity of steel.

OpenStudy (michele_laino):

Hint: The bending work is the work done by the external force. That work will increase the internal energy of the clip paper. So we can write: "Work done by external force = increasing of the internal energy of the clip paper", namely: \[\Large F \times d = c \times M \times \Delta t\] where: F is the external force, d is the distance = 5 cm, M is the mass of the clip paper, c is the specific heat capacity of the steel, and \Delta t is the requested change in temperature. So we have: \[\Large \Delta t = \frac{{Fd}}{{cM}}\] Please make sure to express the distance in meters and the mass of the clip paper in Kilograms

OpenStudy (anonymous):

@Michele_Laino ... dude... shes trying to figure out how much heat is being created at the focal point of the bending paperclip not rocket science. I mean I love physics but man what you said was hard to get into my head! I only kinda got what you were saying. I really think you deserve the medal but you should seriously simplify that down man. Simple easy terms for us simple folk :3

OpenStudy (michele_laino):

I always use simple terms, Lol! @DJ_Timedrop